OpenStudy (maheshmeghwal9):

Given a function f(x). Find at least one root of the equation f(x) = f(a)

mathslover (mathslover):

x=a?

OpenStudy (maheshmeghwal9):

hm yaa:) but how?

mathslover (mathslover):

let f(x) = x+3 so f(x) = f(a) --> given f(a) = a+3 x+3=a+3 x=a

mathslover (mathslover):

I don't know whether I am right or wrong..

OpenStudy (maheshmeghwal9):

oh ok! but is there any use of remainder theorem?

mathslover (mathslover):

yes wait lemme think

OpenStudy (maheshmeghwal9):

ok:)

mathslover (mathslover):

r u sure that i am right with x = a?

OpenStudy (maheshmeghwal9):

yeah this answer is also given in my book:D

mathslover (mathslover):

ok so i got it right so lemme give you explanation with remainder theorem

OpenStudy (maheshmeghwal9):

ok:)

mathslover (mathslover):

I do have remainder = 0 u agree?

mathslover (mathslover):

if i divide f(x) / f(a) then remainder =0

OpenStudy (maheshmeghwal9):

?

OpenStudy (maheshmeghwal9):

sorry i didn't gt u:(

mathslover (mathslover):

given f(x) = f(a) f(x) = f(a) * 1 f(x)/f(a) = 1 this means if we divide f(x) by f(a) the remainder = 0

OpenStudy (anonymous):

@mathslover where did u get 3

OpenStudy (maheshmeghwal9):

how did u get this? divide f(x) / f(a) then remainder =0

mathslover (mathslover):

since f(x) = f(a) what is the remainder when x-3 is divided by x-3 @maheshmeghwal9 ?

OpenStudy (maheshmeghwal9):

0.

mathslover (mathslover):

good so if we dvide f(x) by f(a) we get remainder = 0 right?

mathslover (mathslover):

since f(x)=f(a) hence f(x)/f(a) = f(x)/f(x) => remainder = 0

OpenStudy (maheshmeghwal9):

ya right:)

OpenStudy (maheshmeghwal9):

oh ok i gt it:) thanx @mathslover :D

mathslover (mathslover):

so f(a) is the factor of f(x)

mathslover (mathslover):

that is x-a is a factor of f(x)

mathslover (mathslover):

hence x-a = remainder x-a= 0 x=a proved

OpenStudy (maheshmeghwal9):

:) thanx:)

mathslover (mathslover):

no problem and best of luck

OpenStudy (maheshmeghwal9):

stu:)

mathslover (mathslover):

stu ?

mathslover (mathslover):

what does that mean @maheshmeghwal9

OpenStudy (maheshmeghwal9):

@mathslover sorry for wait reply. actually stu = same to u:D